OpenStudy (goformit100):

Identify the correct x-intercept and y-intercept of the equation 15x - 5y = 30. x-intercept = (-2, 0); y-intercept = (0, 6) x-intercept = (2, 0); y-intercept = (0, -6) x-intercept = (0, -6); y-intercept = (2, 0) x-intercept = (0, 6); y-intercept = (-2, 0)

OpenStudy (goformit100):

@Machida

OpenStudy (anonymous):

y=0, x=2 x=0, y= -6

OpenStudy (anonymous):

why are you asking this @goformit100 ? Lol to raise your score?? HmmM

OpenStudy (rane):

rearrange your equation ist as in y = mx + c

OpenStudy (anonymous):

so, x (0,2) and y (0,-6) I know @goformit100 you just test me lmaooo

OpenStudy (goformit100):

@doulikepiecauseidont does it means if I am 99 i am genius. I am still a OS user.. I can ask my doubt. I don't worry ever about Smart Score.

OpenStudy (anonymous):

Or even easier take the whole equation and divide by 30 and you get all the intercepts at once\[\frac{ x }{ 2 }-\frac{ y }{ 6}=1\]

OpenStudy (goformit100):

I am still a student

OpenStudy (goformit100):

And thanks for the answer @doulikepiecauseidont

OpenStudy (goformit100):

Thankyou @Machida too

OpenStudy (anonymous):

And oh, ok lol Ijs you were asking those project euler questions and i was like :o, where does he get these questions from?

OpenStudy (goformit100):

I get this questions from my teacher and school book.

OpenStudy (goformit100):

*these

OpenStudy (rsadhvika):

give me your teacher number

OpenStudy (anonymous):

No i meant those questions like the sum a 3 digit numbers factorials is that same number, fid the number

OpenStudy (anonymous):

find*

OpenStudy (goformit100):

kyu doon mai number ?

OpenStudy (rsadhvika):

what kind of teacher is she, mixing phd level problems which u have been asking since few days back, and now the kindergarden level problems... Is she alright ?

OpenStudy (anonymous):

Aint it^^